# Hypergames: Modeling Misaligned Perceptions and Nested Beliefs for Multi-agent Systems

### v1

@@ 4 Discussion of Survey Results
Paper
Dynamics
Concept
Formalization
Domain [^1] CS: Cybersecurity, SS: Social Simulations, GT: Game Theory, COM: Communications, ROB: Robotics
Sub-
domain [^2] DD: Defensive Deception, ADM: Attack-Defense Modeling, TOM: Thoery of Mind, CD: Coalition Detection, EOB: Evolution of Beliefs, P-E: Pursuit-Evasion, SC: Semantic Communication, AV: Autonomous Vehicles, CT: Control Theory
Method
Fidelity
Task
[37] Repeated HNF HEU-based CS DD Practical Partial
Reasoning,
Uncertainty
Measure
[23] One-shot MLH L-th Order CS ADM Theoretical Complete Reasoning
[82] Repeated MLH L-th Order CS DD Experimental Complete Reasoning
[72] Repeated MLH Stackelberg CS ADM Experimental Complete Learning
[42] Repeated MLH Graph-based CS ADM Theoretical Complete Reasoning
[69] One-shot MLH
Flattened
L-th Order
SS TOM Practical Complete Reasoning
[47] Repeated MLH Graph-based CS ADM Theoretical Complete Reasoning
[43] One-shot MLH Graph-based CS DD Theoretical Complete Reasoning
[46] One-shot MLH L-th Order SS CD Theoretical Complete Learning
[50] Repeated MLH Graph-based SS ADM Theoretical Complete Planning
[83] Repeated MLH HNE-based GT DC Theoretical Conceptual Reasoning
[28] Repeated MLH Graph-based GT EOB Theoretical Complete Learning
[24] Repeated MLH Graph-based GT EOB Theoretical Complete Learning
[25] Repeated MLH Graph-based GT EOB Theoretical Complete
Learning,
Uncertainty
Measure
[79] Repeated HNF
Stackelberg,
HEU-based
CS DD Practical Partial
Reasoning,
Uncertainty
Measure
[2] Repeated HNF
HEU-based
CS DD Practical Partial
Reasoning
[52] Repeated MLH
Flattened,
L-th Order
GT P-E Experimental Complete
Reasoning,
Learning
[34] Repeated HNF HNF CS ADM Experimental Complete Learning
[71] Repeated MLH Stackelberg COM SC Experimental Complete Learning
[18] Repeated MLH Perceptual ROB AV Practical Conceptual Planning
[48] One-shot MLH L-th Order ROB CT Practical Complete
Uncertainty
Measure
[39] Repeated MLH Graph-based CS DD Experimental Complete Reasoning
[66] Repeated
MLH,
HNF
L-th Order,
HEU-based
CS DD Practical Complete Reasoning
[73] Repeated HNF HEU-based CS ADM Experimental Partial
Reasoning,
Learning
[35] One-shot MLH Stackelberg ROB AV Theoretical Complete Reasoning
@@
Paper
Dynamics
Concept
Formalization
Domain [^1] CS: Cybersecurity, GT: Game Theory, MAS: Multi-agent Systems, ROB: Robotics, SS: Social Simulations
Sub-
domain [^2] ADM: Attack-Defense Modeling, EOB: Evolution of Beliefs, TOM: Thoery of Mind, DD: Defensive Deception, CT: Control theory, AV: Autonomous Vehicles, TD: Trojan Detection, FUZ: Fuzzy Game Theory
Method
Fidelity
Task
[4] Repeated MLH
L-th Order,
HNE-based
CS ADM Theoretical Complete Reasoning
[26] Repeated MLH Graph-based GT EOB Theoretical Complete
Reasoning,
Learning
[1] Repeated MLH
Bayesian
L-th Order
MAS TOM Practical Conceptual
Reasoning,
Learning
[5] One-shot MLH L-th Order CS ADM Theoretical Complete Reasoning
[17] Repeated HNF HEU-based CS DD Experimental Complete
Reasoning,
Uncertainty
Measure
[40] One-shot MLH L-th Order ROB CT Theoretical Complete
Reasoning,
Planning
[49] Repeated MLH Graph-based ROB AV Practical Complete Reasoning
[62] Repeated MLH L-th Order GT SG Experimental Complete Planning
[74] Repeated MLH Graph-based CS DD Theoretical Complete Reasoning
[80] Repeated
MLH,
HNF
L-th Order,
HEU-based
CS DD Practical Complete
Reasoning,
Learning
[14] One-shot MLH
Stackelberg,
HNE-based
CS ADM Theoretical Complete Reasoning
[29] Repeated MLH Graph-based CS ADM Theoretical Complete
Reasoning,
Learning
[27] Repeated MLH Graph-based GT EOB Theoretical Complete
Reasoning,
Learning
[84] One-shot HNF HEU-based CS ADM Practical Partial
Reasoning,
Uncertainty
Measure
[41] One-shot MLH Graph-based CS ADM Theoretical Complete
Reasoning
[70] One-shot MLH
Flattened
L-th Order
SS TOM Experimental Conceptual
Reasoning
[36] One-shot MLH Stackelberg ROB AV Theoretical Complete
Reasoning
[19] Repeated MLH L-th Order CS TD Experimental Complete
Reasoning
[64] One-shot MLH L-th Order GT FUZ Theoretical Complete
Reasoning
@@ 4.1.1 Overview of Model Usage
Figure 9: Distribution of core hypergame concepts implemented per application domain. HNF entails works inspired by Vane’s hypergame normal form [77], Multi-level Hypergames are based on Bennet’s original concept [8] or Wang et al.’s extended formalism [81]. Hybrid approaches use concepts from both main concepts.
@@ 4.2 Agent-based Hypergame Languages
While the extended formalisms by [81] and [77] concretized hypergame frameworks and provided guidelines for hypergame analysis, there is no unified formal language for reasoning about hypergames. Motivated by this gap, hypergame markup language (HML) was developed by [12] to formally encode hypergame models for his hypergame analysis software: HYPANT. HML supports subjective subgames based on player perception but is limited to static utilities and does not support functional utility definitions or dynamic belief updates. Another hypergame analysis software, HAT [31], relies on XML-based structured representation of hypergames. While HAT supports iterative runs and belief updates, it relies on populating Vane’s hypergame normal form representation-based structures. [45] also mention the Security Policy Assistant (SPA) developed by Vane to assist in sensitive decision-making, such as evaluating foreign disclosure of classified information. SPA is not publicly available and we could not find any material for a more detailed analysis.
Despite the aforementioned tools and preliminary modeling languages, an agent-based hypergame modeling language or simulation platform has not emerged. While we encountered various works in our survey that integrate hypergame-theoretic models with linear-temporal-logic-based planning, they do not define a hypergame-specific description language. Future hypergame modeling efforts may take inspiration from alternative frameworks that address information asymmetry and nested beliefs to develop a hypergame-theoretic formal language grounded in logic, such as epistemic logic model that enable dynamic belief updates and planning with iterated knowledge [13, 3], GDL-III providing a formal specification language for general game playing with games of incomplete information [67], or TOMA extending epistemic logic to formalize the theory of mind [21].
@@ 5 Related Work
While Bennett initially developed hypergame theory as an analytical tool for structuring complex strategic scenarios, subsequent research has explored more dynamic applications. [9] demonstrated the use of hypergame theory in real-time decision-making during a shipping crisis. Later, the development of Vane’s HNF provided a more methodological version with potential applications in dynamic, agent-based models. The potential is articulated in his paper addressing decision-theoretic agents [75] and his brief, high-level review of hypergames [76], where the connection to the Observe-Orient-Decide-Act paradigm was also established – a concept well-aligned with human-cognition-inspired agent theories such as the belief-desire-intention (BDI) architecture [68, 65]. Jointly, including the additional formalism provided by [81], while these advancements of hypergame theory offer a richly formal and generalizable framework, up to this point, this line of research remains unconnected to the agent-based and multi-agent systems literature.
@@
[45] presented a broad review of hypergame theory as a framework for modeling conflict, misperception, and deception, particularly in adversarial domains like military planning and cyber defense. Their review emphasizes the theoretical basis of hypergames and highlights how layered beliefs and misaligned perceptions give rise to strategic opportunities unavailable in classical game theory or decision theory. However, their review primarily focused on illustrative, mostly analytical examples rather than methodological analyses.
@@
[85] offered a dual-framework perspective, reviewing defensive deception from both machine learning and game-theoretic viewpoints. Although hypergame theory is featured in their survey through selected analytical and practical examples, the coverage remains limited and application-specific, without substantial linkage to broader multi-agent systems research. Complementing this work, [38] presented a systematic review of game-theoretic defense models against advanced persistent threats. While not dedicated to hypergames and, as such, lacking formal descriptions of hypergame models, their review acknowledges the value of hypergame theory in modeling adversarial perturbations and covers three additional hypergame-based works on defensive deception. Crook’s doctoral dissertation [16] represents a unique contribution bridging computability theory and hypergames, involving an exploration of how attack graphs in cybersecurity can be recast as hypergames. While preliminary, this work proposes several directions for formalizing adaptive, layered-belief hypergames in cybersecurity with links to computable analysis. Finally, [22] addressed hypergame theory in their review of game-theoretic models in cyber deception. While their discussion is brief and only on a high level, the paper explicitly recognizes that attacker–defender interactions often exhibit asymmetries traditional games cannot capture. The consideration of hypergames for the adaptability of defender strategies progresses the practical generalizability of hypergame theory.
@@ 6 Conclusions
Despite early efforts such as HML [12] and HAT [31], there remains no formal language or simulation framework explicitly designed for modeling agent-based hypergames. Given the identified lack of a unified representational formalism and the need for structured support for recursive belief modeling and dynamic perception updates, future work should prioritize the development of dedicated hypergame description languages tailored to agent-based applications. Inspiration may be drawn from planning frameworks rooted in epistemic logic [13, 3], general game-playing languages like GDL-III [67], or formal ToM modeling tools such as TOMA [21].
Across the surveyed domains, hypergames have been successfully employed to model nested beliefs, misaligned perceptions, and deception; however, few works explicitly link these models to established cognitive agent architectures. As recursive reasoning becomes increasingly central in agent design, integrating hypergame theory with human-inspired frameworks such as BDI [68, 65] could enrich symbolic reasoning and offer interpretable mechanisms for belief formation and adaptation in dynamic environments. Furthermore, while we encountered only a handful of hypergame-based approaches addressing large language models (LLMs), the growing prevalence of human–agent [60] and agent–agent misalignment [44], along with the application of theory of mind in agentic systems [11, 59, 21], suggests promising opportunities for hypergame theory to formalize ToM-like models capable of capturing and mitigating agent misalignment.

### v2

@@ 4 Discussion of Survey Results
Paper Dynamics Concept Formalization Domain [^1] CS: Cybersecurity, SS: Social Simulations, GT: Game Theory, COM: Communications, ROB: Robotics Sub- domain [^2] DD: Defensive Deception, ADM: Attack-Defense Modeling, TOM: Thoery of Mind, CD: Coalition Detection, EOB: Evolution of Beliefs, P-E: Pursuit-Evasion, SC: Semantic Communication, AV: Autonomous Vehicles, CT: Control Theory Method Fidelity Task
[jia2024defense] Repeated HNF HEU-based CS $D$ Practical Partial Reasoning, Uncertainty Measure
[guiterrez2018] One-shot MLH L-th Order CS ADM Theoretical Complete Reasoning
[bowei2020iobt] Repeated MLH L-th Order CS $D$ Experimental Complete Reasoning
[TANG2024103871] Repeated MLH Stackelberg CS ADM Experimental Complete Learning
[kulkarni2021hypergamegraphs] Repeated MLH Graph-based CS ADM Theoretical Complete Reasoning
[trencsenyi2025approximatinghumanstrategicreasoning] One-shot MLH Flattened L-th Order $S$ TOM Practical Complete Reasoning
[kulkarni2021labeling] Repeated MLH Graph-based CS ADM Theoretical Complete Reasoning
[kulkarni2020decoyallocation] One-shot MLH Graph-based CS $D$ Theoretical Complete Reasoning
[kulkarni2025coalitiondetection] One-shot MLH L-th Order $S$ CD Theoretical Complete Learning
[lening2023dynamichypergames] Repeated MLH Graph-based $S$ ADM Theoretical Complete Planning
[yuan2024equilibrium] Repeated MLH HNE-based GT DC Theoretical Conceptual Reasoning
[gharesifard2012evolution] Repeated MLH Graph-based GT EOB Theoretical Complete Learning
[gharesifard2010evolution] Repeated MLH Graph-based GT EOB Theoretical Complete Learning
[gharesifard2011exploration] Repeated MLH Graph-based GT EOB Theoretical Complete Learning, Uncertainty Measure
[wan2021foureyedefensivedeceptionbased] Repeated HNF Stackelberg, HEU-based CS $D$ Practical Partial Reasoning, Uncertainty Measure
[anwar2020] Repeated HNF HEU-based CS $D$ Practical Partial Reasoning
[li2024hypergameQlearning] Repeated MLH Flattened, L-th Order GT P-E Experimental Complete Reasoning, Learning
[house2010hypergame] Repeated HNF HNF CS ADM Experimental Complete Learning
[thomas2024hypergametheorydecentralizedresource] Repeated MLH Stackelberg COM SC Experimental Complete Learning
[dharmadhikari2021dronesexploring] Repeated MLH Perceptual ROB AV Practical Conceptual Planning
[kahn2022traffic] One-shot MLH L-th Order ROB CT Practical Complete Uncertainty Measure
[kulkarni2024integratedresourceallocationstrategy] Repeated MLH Graph-based CS $D$ Experimental Complete Reasoning
[shen2025iot] Repeated MLH, HNF L-th Order, HEU-based CS $D$ Practical Complete Reasoning
[thukkaraju2023mia] Repeated HNF HEU-based CS ADM Experimental Partial Reasoning, Learning
[he2022leader] One-shot MLH Stackelberg ROB AV Theoretical Complete Reasoning
@@
Paper Dynamics Concept Formalization Domain [^1] CS: Cybersecurity, GT: Game Theory, MAS: Multi-agent Systems, ROB: Robotics, SS: Social Simulations Sub- domain [^2] ADM: Attack-Defense Modeling, EOB: Evolution of Beliefs, TOM: Thoery of Mind, DD: Defensive Deception, CT: Control theory, AV: Autonomous Vehicles, TD: Trojan Detection, FUZ: Fuzzy Game Theory Method Fidelity Task
[bakker2020infomanipulation] Repeated MLH L-th Order, HNE-based CS ADM Theoretical Complete Reasoning
[gharesifard2011learningequilibria] Repeated MLH Graph-based GT EOB Theoretical Complete Reasoning, Learning
[aitchison2021hiddenrolegames] Repeated MLH Bayesian L-th Order MAS TOM Practical Conceptual Reasoning, Learning
[bakker2021metagames] One-shot MLH L-th Order CS ADM Theoretical Complete Reasoning
[Cho2019deceptiongames] Repeated HNF HEU-based CS $D$ Experimental Complete Reasoning, Uncertainty Measure
[kulkarni2019reactivegames] One-shot MLH L-th Order ROB CT Theoretical Complete Reasoning, Planning
[lv2024synthesisforUAV] Repeated MLH Graph-based ROB AV Practical Complete Reasoning
[chongyang2023stochasticgames] Repeated MLH L-th Order GT SG Experimental Complete Planning
[udupa2024reactivesynthesissensorrevealing] Repeated MLH Graph-based CS $D$ Theoretical Complete Reasoning
[wan2023APT] Repeated MLH, HNF L-th Order, HEU-based CS $D$ Practical Complete Reasoning, Learning
[cheng2022stackelberg] One-shot MLH Stackelberg, HNE-based CS ADM Theoretical Complete Reasoning
[gharesifard2014stealthydeception] Repeated MLH Graph-based CS ADM Theoretical Complete Reasoning, Learning
[gharesifard2011stealthy] Repeated MLH Graph-based GT EOB Theoretical Complete Reasoning, Learning
[yoon2025mia] One-shot HNF HEU-based CS ADM Practical Partial Reasoning, Uncertainty Measure
[kulkarni2021reachabilitygames] One-shot MLH Graph-based CS ADM Theoretical Complete Reasoning
[trencsenyi2025influencehumaninspiredagenticsophistication] One-shot MLH Flattened L-th Order $S$ TOM Experimental Conceptual Reasoning
[hai2023uavcommunication] One-shot MLH Stackelberg ROB AV Theoretical Complete Reasoning
[tapadhir2020] Repeated MLH L-th Order CS TD Experimental Complete Reasoning
[song2009fuzzyhypergame] One-shot MLH L-th Order GT FUZ Theoretical Complete Reasoning
@@ 4.1.1 Overview of Model Usage
Figure 9: Distribution of core hypergame concepts implemented per application domain. HNF entails works inspired by Vane’s hypergame normal form [vane2000thesis], Multi-level Hypergames are based on Bennet’s original concept [Bennett1980] or Wang et al.’s extended formalism [Wang1988]. Hybrid approaches use concepts from both main concepts.
@@ 4.2 Agent-based Hypergame Languages
While the extended formalisms by [Wang1988] and [vane2000thesis] concretized hypergame frameworks and provided guidelines for hypergame analysis, there is no unified formal language for reasoning about hypergames. Motivated by this gap, hypergame markup language (HML) was developed by [brumley2003hypant] to formally encode hypergame models for his hypergame analysis software: HYPANT. HML supports subjective subgames based on player perception but is limited to static utilities and does not support functional utility definitions or dynamic belief updates. Another hypergame analysis software, HAT [Gibson2013hypergame], relies on XML-based structured representation of hypergames. While HAT supports iterative runs and belief updates, it relies on populating Vane’s hypergame normal form representation-based structures. [kovach2015hypergamereview] also mention the Security Policy Assistant (SPA) developed by Vane to assist in sensitive decision-making, such as evaluating foreign disclosure of classified information. SPA is not publicly available and we could not find any material for a more detailed analysis.
Despite the aforementioned tools and preliminary modeling languages, an agent-based hypergame modeling language or simulation platform has not emerged. While we encountered various works in our survey that integrate hypergame-theoretic models with linear-temporal-logic-based planning, they do not define a hypergame-specific description language. Future hypergame modeling efforts may take inspiration from alternative frameworks that address information asymmetry and nested beliefs to develop a hypergame-theoretic formal language grounded in logic, such as epistemic logic model that enable dynamic belief updates and planning with iterated knowledge [baltag2008qualitative, bolander2011epistemic], GDL-III providing a formal specification language for general game playing with games of incomplete information [thielscher2017gdl3], or TOMA extending epistemic logic to formalize the theory of mind [erdogan2025toma].
@@ 5 Related Work
While Bennett initially developed hypergame theory as an analytical tool for structuring complex strategic scenarios, subsequent research has explored more dynamic applications. [BENNETT1981shipping] demonstrated the use of hypergame theory in real-time decision-making during a shipping crisis. Later, the development of Vane’s HNF provided a more methodological version with potential applications in dynamic, agent-based models. The potential is articulated in his paper addressing decision-theoretic agents [vane2000dtgt] and his brief, high-level review of hypergames [vane2006advances], where the connection to the Observe-Orient-Decide-Act paradigm was also established – a concept well-aligned with human-cognition-inspired agent theories such as the belief-desire-intention (BDI) architecture [TWEEDALE2007bdi, Simari2011bdi]. Jointly, including the additional formalism provided by [Wang1988], while these advancements of hypergame theory offer a richly formal and generalizable framework, up to this point, this line of research remains unconnected to the agent-based and multi-agent systems literature.
@@
[kovach2015hypergamereview] presented a broad review of hypergame theory as a framework for modeling conflict, misperception, and deception, particularly in adversarial domains like military planning and cyber defense. Their review emphasizes the theoretical basis of hypergames and highlights how layered beliefs and misaligned perceptions give rise to strategic opportunities unavailable in classical game theory or decision theory. However, their review primarily focused on illustrative, mostly analytical examples rather than methodological analyses.
@@
[zhu2021survey] offered a dual-framework perspective, reviewing defensive deception from both machine learning and game-theoretic viewpoints. Although hypergame theory is featured in their survey through selected analytical and practical examples, the coverage remains limited and application-specific, without substantial linkage to broader multi-agent systems research. Complementing this work, [khalid2023gtAPT] presented a systematic review of game-theoretic defense models against advanced persistent threats. While not dedicated to hypergames and, as such, lacking formal descriptions of hypergame models, their review acknowledges the value of hypergame theory in modeling adversarial perturbations and covers three additional hypergame-based works on defensive deception. Crook’s doctoral dissertation [crook2024thesis] represents a unique contribution bridging computability theory and hypergames, involving an exploration of how attack graphs in cybersecurity can be recast as hypergames. While preliminary, this work proposes several directions for formalizing adaptive, layered-belief hypergames in cybersecurity with links to computable analysis. Finally, [Fugate_Ferguson-Walter_2019] addressed hypergame theory in their review of game-theoretic models in cyber deception. While their discussion is brief and only on a high level, the paper explicitly recognizes that attacker–defender interactions often exhibit asymmetries traditional games cannot capture. The consideration of hypergames for the adaptability of defender strategies progresses the practical generalizability of hypergame theory.
@@ 6 Conclusions
Despite early efforts such as HML [brumley2003hypant] and HAT [Gibson2013hypergame], there remains no formal language or simulation framework explicitly designed for modeling agent-based hypergames. Given the identified lack of a unified representational formalism and the need for structured support for recursive belief modeling and dynamic perception updates, future work should prioritize the development of dedicated hypergame description languages tailored to agent-based applications. Inspiration may be drawn from planning frameworks rooted in epistemic logic [baltag2008qualitative, bolander2011epistemic], general game-playing languages like GDL-III [thielscher2017gdl3], or formal ToM modeling tools such as TOMA [erdogan2025toma].
Across the surveyed domains, hypergames have been successfully employed to model nested beliefs, misaligned perceptions, and deception; however, few works explicitly link these models to established cognitive agent architectures. As recursive reasoning becomes increasingly central in agent design, integrating hypergame theory with human-inspired frameworks such as BDI [TWEEDALE2007bdi, Simari2011bdi] could enrich symbolic reasoning and offer interpretable mechanisms for belief formation and adaptation in dynamic environments. Furthermore, while we encountered only a handful of hypergame-based approaches addressing large language models (LLMs), the growing prevalence of human–agent [strachan2024testing] and agent–agent misalignment [Kierans2025misalignment], along with the application of theory of mind in agentic systems [bosse2011recursive, rocha2023tomReview, erdogan2025toma], suggests promising opportunities for hypergame theory to formalize ToM-like models capable of capturing and mitigating agent misalignment.